Search This Blog


Sunday 28th December 2008 ~ Gemini Gremlins

Problems with the B9s seem to be mounting up. Today 37572 was still out of traffic with its CCTV not working and it had been joined at Rowan road by 37570 which has a gearbox problem. The serviceable B7 contingent is now down to one with 20352 still at Volvo (Ely) and 20353 stopped at Lynn with its screen unserviceable. Out on the road today were Geminis 37563/5/7/8/9, 37571/7. The other two turns were 20351 (19.59 Lynn - Peterborough) and 20105 which did the 20.59 westbound. B10Ms at Lynn tonight were 20103/9/21/31 plus out of use 20120/141. 20121 doesn't appear to have worked since before Christmas as it still has a Hunstanton schools contract number on the blinds. What happened to 20111 today is unknown - may be it was out at Hunstanton this evening ?

No comments: